# Minimalist Market Design: A Framework for Economists with Policy Aspirations

### v3

@@ 2.2 Three Pillars of Minimalist Market Design
To address these challenges, minimalist market design is built upon three foundational pillars:
@@
    Identify the Mission. Decision makers and stakeholders across various economic, social, and political environments may be flexible about the technical aspects of an institution—such as its allocation mechanisms or operational rules—but they often hold firm views on the underlying policy objectives. Thoroughly understanding these objectives is crucial for effective design.
@@
    Avoid Nonessential Changes. As an outsider critic, you may not fully understand the role of every aspect of an institution. Therefore, it is prudent to avoid altering elements that do not affect the primary objectives, minimizing the risk of unintended consequences.
@@
    Incorporate a Persuasion Strategy. Your ideas, as an outsider, are likely to face resistance from decision makers. To overcome this, it is essential to build a compelling case to persuade them to undertake potentially costly reforms.
Based on these pillars, minimalist design economists addressing the failures of a flawed institution first identify the key policy objectives, determine the root causes of the failures, and then address them by directly targeting their sources. To maximize practical relevance and persuade stakeholders of the value of these ideas, they typically develop custom models tailored to specific applications and create novel theories to identify and resolve the root causes of failure.
@@ 2.2.1 Establishing the Institutional Mission
With this clarification on terminology, minimalist market design is conceived as a framework that serves legitimate objectives, [^6] In principle, the existing institution may embed various illegitimate objectives. By maintaining as much of the existing institution as possible, a reform guided by minimalist market design may also inherit some of these aspects. In applications where the legitimate objectives imply a single design, this cannot happen. However, in applications where multiple designs satisfy the legitimate objectives, it can. Therefore, the best practice in those cases is to provide a complete description of these institutions, including institutions where the illegitimate objectives are weeded out. This exercise is referred to as a “full characterization” in axiomatic methodology, an analytical framework that often plays a key role in minimalist market design. and is largely intended for institutions governed by honest and democratically elected officials who genuinely aim to enhance these institutions. [^7] Although the framework could be misused to serve the selfish objectives of a politically motivated or malicious decision maker or interest group, I would not consider such misuse as an exercise in minimalist market design, since the framework was developed for public service.
@@
An illustrative example of this can be found in Section 6.12.1. In 2007, a strategy-proof mechanism was mandated across England for the allocation of public school seats through a reform of the Education Code. The primary justification given by Alan Johnson, the Education Secretary at the time, was that the previous mechanism “forced many parents to play an ‘admissions game’ with their children’s future.” Thus, strategy-proofness was an objective of the policymakers not solely for its consequences on the outcomes but rather because of the values it promoted.
@@
Assisting policymakers in formulating and accommodating their normative goals has been a productive avenue for minimalist design economists. As noted by Li (2017), while policymakers may possess deep knowledge of their specific environments, they often struggle to articulate their ethical requirements with precision. Economists, with their comparative advantage in defining the normatively relevant properties of complex systems, can help translate these ethical considerations into formal concepts or mechanisms.
@@ 2.2.2 Avoiding Nonessential Changes
Although a design economist may not uncover every policy goal of stakeholders, existing institutions often contain critical information about the system that is difficult, if not impossible, to obtain otherwise. Even when an institution is deeply flawed, it may still reflect important aspects of the problem that an aspiring market designer might overlook. In this context, the second pillar of minimalist market design aligns closely with Chesterton’s Fence, a well-known principle articulated by G.K. Chesterton in the opening paragraph of the chapter “The Drift from Domesticity” in his book The Thing (Chesterton, 1929):
@@
In minimalist market design, Chesterton’s Fence is reflected in the following operational principle: when proposing a reform, only alter elements that are “clearly broken” with respect to stakeholders’ objectives—the “root causes” of failures, which often stem from a lack of formal expertise. Beyond these essential changes, refrain from making others. Consistent with Chesterton’s main message, this approach helps prevent unintended consequences by avoiding the disruption of essential institutional elements. Moreover, it presents stakeholders with designs that closely resemble the original institution, minus its shortcomings. Ideally, the central planner should feel, “this is the institution we intended to design in the first place,” thereby supporting the third pillar of minimalist market design as well.
A prevailing view in market design suggests that, all else being equal—such as the potential value generated by the intervention—the closer a proposed reform aligns with the existing system, the more likely it is to gain stakeholder support. However, no other institution design framework places as much emphasis on avoiding changes not directly related to institutional failures, largely due to the constraints of their underlying technical frameworks. This steadfast pursuit of minimalism is one of the most distinctive features of minimalist market design, and it is precisely what gives the framework its name.
@@ 2.2.3 Forging a Design Framework with a Built-in Persuasion Strategy
Building on the first two pillars—identifying the mission and avoiding nonessential changes—the third pillar highlights the crucial role of persuasion in driving reforms. Influencing real-world policies or prompting costly institutional changes from the outside can be challenging, even with exceptional research, as decision makers often resist external critiques due to vested interests in maintaining the status quo.
One of the primary motivations for minimalist market design is to address these challenges when a design economist functions as an external critic rather than a commissioned guide. In such contexts, the need for change is rarely acknowledged, and a critical perspective from an eager researcher is often met with resistance.
To overcome this resistance, the research model must be highly realistic, capturing all essential aspects of the problem—especially when addressing decision makers who are reluctant to change. In my experience, this often necessitates developing a custom-made theory rather than relying on simplified, off-the-shelf research. Unlike experts who are commissioned to guide reforms—often brought in to address urgent crises—an aspiring design economist cannot depend on simplified analyses to succeed, particularly when attempting to dismantle an existing institution. Therefore, for such ambitions to bear fruit, the model must be meticulously realistic, and the arguments for reform must be compelling and airtight.
Consistent with the first two pillars of minimalist market design, once a realistic model is crafted, the proposed solutions must closely align with the aims of policymakers and other stakeholders, ideally without inadvertently compromising any unidentified objectives. Essentially, it must be demonstrated that the proposed design better meets the stakeholders’ goals than the current system. Such reforms provide genuine value to decision makers and are easier to justify to superiors or constituents.
@@
Within this framework, the market designer’s initial role is that of a supportive partner. The only interference with the existing institution involves correcting or enhancing flawed elements that require formal expertise. This approach helps establish trust and allows stakeholders to ”own” the proposed design. By adhering to this framework, design economists avoid adopting any normative positions—intentional or unintentional—beyond aligning with the stakeholders’ own objectives during initial interactions with likely skeptical parties. This strategy helps stakeholders perceive that the proposed reform has no ideological agenda or ulterior motive. Providing pro bono support can also be highly valuable—perhaps even necessary—in building trust.
@@ 2.3 Characteristics and Features of Minimalist Market Design
  • Mission-Driven Focus: Centers on the objectives of decision makers, system operators, and other stakeholders, extending beyond the typical focus of mainstream economists. Frequently derives these objectives from historical contexts.
@@ 3 From Theory to Practice: The Role of Innovation Models
According to common wisdom, market design is considered a field that builds on mechanism design (Hurwicz, 1960, 1972, 1973) as its main theoretical foundation, but it places more emphasis on institutional details and aims to provide policy-relevant insights. Throughout the monograph, I refer to the spectrum of research that encompasses both foundational research in mechanism design and application-oriented market design as design economics. [^11] This spectrum also includes research in applied mechanism design, which aims not to inform policy within a specific institutional setting but to understand properties that hold across a variety of settings, provided they meet certain technical assumptions.
@@
“A lesson from this experience of theorists in policymaking is that the real value of the theory is in developing intuition. [$\ldots$] What the theorists found to be the most useful in designing the auction and advising the bidders was not complicated models that try to capture a lot of reality at the cost of relying on special functional forms.”
@@
“Research, which we understand as an unscheduled quest for new knowledge and the creation of new inventions, whose outcome cannot be predicted in advance [$\cdots$] Development, which is a scheduled activity with a well-defined outcome in a specified time frame, aimed at the marketplace.” (p. 11)
@@
The indirect relationship between theory and practice in the engineering approach to market design is largely shaped by the demands of commissioned projects. Design economists in such settings lack the flexibility to choose their problems or timelines; they must deliver practical solutions under strict deadlines. As Narayanamurti and Odumosu (2016) note, the pressures of completing a “scheduled” activity often leave little room for advancing theoretical boundaries. This dynamic, however, is not unique to engineering approach to market design—it reflects a broader paradigm articulated by Bush (1945), which envisions a linear, unidirectional relationship between scientific discovery and technological innovation. This model has profoundly influenced U.S. science and technology policy since World War II.
In contrast to the pioneers of market design who were engaged in commissioned projects, I—as an aspiring design economist—had two significant flexibilities unavailable to them in the late 1990s when I ventured into the world of policy design. I could actively seek out real-world applications with a stronger role for novel theory, and I had ample time to pursue these projects as an “unscheduled quest for new knowledge and the creation of new inventions.”
Freed from these constraints, minimalist market design emerged as an exception to the broader trend shaped by Bush’s linear model. It embraces a holistic approach that integrates research and practice, aligning instead with alternative innovation frameworks advocated by Stokes (1997) and Narayanamurti and Odumosu (2016). To clarify these distinctions, we will first review Bush’s post-war science and technology framework in Section 3.1, then examine Stokes’s Pasteur’s Quadrant model in Section 3.2, and finally discuss the Discovery-Invention Cycles model introduced by Narayanamurti and Odumosu in Section 3.3.
@@ 3.1 Science, the Endless Frontier: Vannevar Bush’s Linear Model of Innovation
Dichotomy Between Basic and Applied Research: Bush argued that “basic research is performed without thought of practical ends.” He suggested that basic research, aiming to broaden knowledge without concern for practical applications, and applied research, seeking to solve specific practical problems, should remain distinct. According to Bush, focusing on one detracts from the other, as “applied research invariably drives out pure research.”
Linearity of the Four Phases of Innovation: Bush further asserted that “basic research is the pacemaker of technological progress.” In his model of innovation, basic research expands knowledge without a specific application in mind. Applied research then uses this knowledge to create practical models and innovations. Development follows, converting research findings into useful outputs like materials, devices, systems, and procedures. Finally, these outputs are put into practical use. Each stage is successive and depends on the previous one.
@@
The fifty years following Bush’s report witnessed unprecedented research advances, during which the U.S. emerged as a global leader in science. Bush’s linear model of innovation, as well as its influence on shaping institutions like the NSF and other government funding bodies, went largely unchallenged until the mid-1990s. This historical legacy, coupled with the funding structures of U.S. federal agencies like the NSF during that period, contributed to the prominence of Bush’s influential theses in many fields, including design economics.
@@ 3.2 Beyond the Linear Model: Pasteur’s Quadrant
“Science and technology policy would appear to be in a state of crisis. There are many indicators that a crisis does exist in the partnership between the federal government and the American research universities, that the terms of the partnership are increasingly being questioned and re-examined. [$\cdots$]
Even if Vannevar Bush is as much a symbol of this period of dramatic progress as its putative architect, it seems fitting that in the 50th anniversary year of the publication of Science: The Endless Frontier, we celebrate the work and the period: reflect on the origins of the Bush paradigm for scientific excellence, take stock of where we currently are in the relationship between science and government, consider whether we are in a period of crisis, and do some serious work on the future shape of the national system of scientific and technological innovation. [$\cdots$]
@@
To address the limitations of Bush’s unidimensional approach, Stokes introduces a framework that incorporates a second dimension: the desire to solve practical problems alongside the quest to expand fundamental understanding. His model divides research into three primary quadrants, [^14] There is also a fourth quadrant where research neither extends fundamental understanding nor serves a practical purpose. each represented by a key historical figure:
@@
Since its introduction in the late 1990s, Stokes’s framework has profoundly influenced innovation policy in the U.S. For example, it inspired the NSF to incorporate a “broader impacts” criterion into its merit review process for grant applications in the early 2010s. Its influence has been so significant that, in 2014, in response to those advocating otherwise, the NSF Assistant Director for Mathematical and Physical Sciences emphasized that the NSF should not entirely abandon support for pure basic research in Bohr’s Quadrant by forcing all basic research into Pasteur’s Quadrant (Crim, 2014):
@@
Stokes’s framework has also influenced innovation policy beyond the U.S. For instance, since 2011, researchers submitting proposals to the Swiss National Science Foundation have been able to designate their applications as “use-inspired.”
To better understand how these quadrants manifest in the field of design economics, let us explore specific examples that illustrate research practices within each quadrant. We will focus on seminal works that have not only advanced the field of design economics but have also contributed to their authors receiving the Nobel Memorial Prize in Economic Sciences, thereby underscoring their significance. We begin with Bohr’s Quadrant.

### v4

@@ 2.2 Three Pillars of Minimalist Market Design
To address these challenges, minimalist market design is built on three foundational pillars, illustrated in Figure 2:
Figure 2: Pillars of Minimalist Market Design.
@@
    Identify the Mission. Decision-makers and other stakeholders across economic, social, and political environments may be flexible about the technical aspects of an institution—such as its allocation mechanisms or operational rules—but they often hold firm views about the underlying policy objectives. Thoroughly understanding these objectives is crucial for effective design.
@@
    Avoid Nonessential Changes. As an outside critic, you may not fully understand the role of every aspect of an institution. Therefore, avoid altering elements that do not affect the primary objectives; doing so minimizes the risk of unintended consequences.
@@
    Incorporate a Persuasion Strategy. As an outsider, your ideas are likely to face resistance from decision-makers. To overcome this, it is essential to build a compelling case to persuade them to undertake potentially costly reforms.
Based on these pillars, minimalist economic designers confronting a flawed institution first identify the key policy objectives, then determine the root causes of failure, and finally address them by targeting those causes directly. To maximize practical relevance and to persuade stakeholders, they typically develop custom, application-specific models and advance new theory to help identify and resolve those root causes.
@@ 2.2.1 Establishing the Institutional Mission
With this clarification on terminology, minimalist market design is conceived as a framework that serves legitimate objectives, [^9] In principle, the existing institution may embed various illegitimate objectives. By maintaining as much of the existing institution as possible, a reform guided by minimalist market design may also inherit some of these aspects. In applications where the legitimate objectives imply a single design, this cannot happen. However, in applications where multiple designs satisfy the legitimate objectives, it can. Therefore, the best practice in those cases is to provide a complete description of these institutions, including institutions where the illegitimate objectives are weeded out. This exercise is referred to as a “full characterization” in axiomatic methodology, an analytical framework that often plays a key role in minimalist market design. and is largely intended for institutions governed by honest and democratically elected officials who genuinely aim to enhance these institutions. [^10] Although the framework could be misused to serve the selfish objectives of a politically motivated or malicious decision-maker or interest group, I would not consider such misuse as an exercise in minimalist market design, since the framework was developed for public service.
@@
An illustrative example of this can be found in Section 6.12.1. In 2007, a strategy-proof mechanism was mandated across England for the allocation of state school seats through a reform of the Education Code. According to Alan Johnson, the Education Secretary at the time, the motivation was that the previous mechanism compelled many parents to engage in strategic behavior in order to secure school places for their children. Thus, strategy-proofness was an objective of the policymakers not solely for its consequences on the outcomes but rather because of the values it promoted.
@@
Assisting policymakers in formulating and accommodating their normative goals has been a productive avenue for minimalist economic designers. As noted by Li (2017), while policymakers may possess deep knowledge of their specific environments, they often struggle to articulate their ethical requirements with precision. Economists, with their comparative advantage in defining the normatively relevant properties of complex systems, can help translate these ethical considerations into formal concepts or mechanisms.
@@ 2.2.2 Avoiding Nonessential Changes
Although an economic designer may not uncover every policy goal of stakeholders, existing institutions often contain critical information about the system that is difficult, if not impossible, to obtain otherwise. Even when an institution is deeply flawed, it may still reflect important aspects of the problem that an aspiring market designer might overlook. In this context, the second pillar of minimalist market design aligns closely with Chesterton’s Fence, a well-known principle articulated by G.K. Chesterton in the opening paragraph of the chapter “The Drift from Domesticity” in his book The Thing (Chesterton, 1929):
@@
In minimalist market design, Chesterton’s Fence is reflected in the following operational principle: when proposing a reform, only alter elements that are “clearly broken” with respect to stakeholders’ objectives—namely, the root causes of failure, which often stem from a lack of formal expertise. Beyond these essential changes, refrain from making any others. Consistent with Chesterton’s main message, this approach helps prevent unintended consequences by avoiding the disruption of essential institutional elements. Moreover, it presents stakeholders with designs that closely resemble the original institution, minus its shortcomings. Ideally, the central planner should feel, “this is the institution we intended to design in the first place,” thereby supporting the third pillar of minimalist market design as well.
This caution has deep antecedents in political philosophy and policy analysis. Edmund Burke’s defense of gradual reform stressed that long-standing institutions embody tacit knowledge accumulated over generations; well-meaning attempts at a full-scale overhaul risk destroying functions not yet understood (Burke, 1790). For Burke, legitimate improvement is piecemeal and evolutionary: it involves correcting faults while preserving the inherited structure, and placing the burden of proof on those who would overturn established arrangements. Together with Chesterton’s Fence, Burkean conservatism provides a normative rationale for changing only what is demonstrably defective—fix what is broken, and respect the rest (Burke, 1790).
Charles Lindblom’s classic account of policymaking as “muddling through” provides a pragmatic complement to this ethos. According to Lindblom, rather than comprehensive, blueprint-style redesign, effective policy change typically proceeds by successive limited modifications—small moves from the status quo that target identifiable problems, invite feedback, and minimize unforeseen spillovers (Lindblom, 1959). Incrementalism is not timidity but a method for learning under complexity and political constraint. In practice, it directs analytic attention toward feasible amendments while preserving institutional components that are functioning tolerably well. That is precisely the minimalist approach: target the root cause, avoid unnecessary interference.
In public finance, Martin Feldstein’s theory of tax reform exemplifies how targeted adjustments can deliver durable change without uprooting foundational structures. Feldstein draws a sharp distinction between designing a system de novo and reforming an existing one, urging policymakers to treat the status quo as the baseline and focus on correcting clear deficiencies rather than pursuing an idealized design (Feldstein, 1976). His approach—identifying concrete failures in the status quo, devising analytically well-grounded and administratively feasible corrections, and implementing these targeted fixes while leaving the rest of the system intact—embodies the minimalist ethos of repairing what is demonstrably broken and nothing more.
Taken together, these perspectives reinforce the prudence of avoiding nonessential changes. Burke frames the normative foundation of inherited wisdom; Lindblom offers an operational logic for incremental correction under complexity; and Feldstein clarifies the distinction between reform and unrestricted design from scratch. Each points to the same rule of thumb that guides minimalist market design: alter only those components that compromise the mission, and preserve the rest, thereby reducing unintended consequences and easing persuasion.
This rule of thumb is not entirely new to the scholarship and practice of market design. A prevailing view holds that all else equal—such as the expected value generated by the intervention—the closer a proposed reform aligns with the existing system, the more likely it is to gain stakeholder support. What distinguishes minimalist market design is the degree to which it operationalizes this insight—elevating the avoidance of nonessential changes from a rule of thumb to a design constraint. No other economic design framework places as much emphasis on avoiding changes not directly related to institutional failures, and this steadfast discipline is precisely what gives the minimalist approach its name.
@@ 2.2.3 Forging a Design Framework with a Built-in Persuasion Strategy
Building on the first two pillars—identifying the mission and avoiding nonessential changes—the third pillar highlights the crucial role of persuasion in driving reforms. Influencing real-world policies or prompting costly institutional changes from the outside can be challenging, even with exceptional research, as decision-makers often resist external critiques due to vested interests in maintaining the status quo.
One of the primary motivations for minimalist market design is to address these challenges when an economic designer functions as an external critic rather than a commissioned guide. In such contexts, the need for change is rarely acknowledged, and a critical perspective from an eager researcher is often met with resistance.
To overcome this resistance, the research model must be highly realistic, capturing all essential aspects of the problem—especially when addressing decision-makers who are reluctant to change. In my experience, this often necessitates developing a custom-made theory rather than relying on simplified, off-the-shelf research. Unlike experts who are commissioned to guide reforms—often brought in to address urgent crises—an aspiring economic designer cannot depend on simplified analyses to succeed, particularly when attempting to dismantle an existing institution. Therefore, for such ambitions to bear fruit, the model must be meticulously realistic, and the arguments for reform must be compelling and airtight.
Consistent with the first two pillars of minimalist market design, once a realistic model is crafted, the proposed solutions must closely align with the aims of policymakers and other stakeholders, ideally without inadvertently compromising any unidentified objectives. Essentially, it must be demonstrated that the proposed design better meets the stakeholders’ goals than the current system. Such reforms provide genuine value to decision-makers and are easier to justify to superiors or constituents.
@@
Within this framework, the market designer’s initial role is that of a supportive partner. The only interference with the existing institution involves correcting or enhancing flawed elements that require formal expertise. This approach helps establish trust and allows stakeholders to “own” the proposed design. By adhering to this framework, economic designers avoid adopting any normative positions—intentional or unintentional—beyond aligning with the stakeholders’ own objectives during initial interactions with likely skeptical parties. This strategy helps convince stakeholders that the proposer has no ideological agenda or ulterior motive. Providing pro bono support can also be highly valuable—perhaps even necessary—in building trust.
@@ 2.3 Characteristics and Features of Minimalist Market Design
  • Mission-Driven Focus: Centers on the objectives of decision-makers, system operators, and other stakeholders, extending beyond the typical focus of mainstream economists. Frequently derives these objectives from historical contexts.
@@ 3 From Theory to Practice: The Role of Innovation Models
According to common wisdom, market design is considered a field that builds on mechanism design (Hurwicz, 1960, 1972, 1973) as its main theoretical foundation, but it places more emphasis on institutional details and aims to provide policy-relevant insights. Throughout the monograph, I refer to the spectrum of research that encompasses both foundational research in mechanism design and application-oriented market design as economic design. [^14] This spectrum also includes research in applied mechanism design, which aims not to inform policy within a specific institutional setting but to understand properties that hold across a variety of settings, provided they meet certain technical assumptions.
@@
“A lesson from this experience of theorists in policymaking is that the real value of the theory is in developing intuition. $\ldots$] What the theorists found to be the most useful in designing the auction and advising the bidders was not complicated models that try to capture a lot of reality at the cost of relying on special functional forms.”
@@
“Research, which we understand as an unscheduled quest for new knowledge and the creation of new inventions, whose outcome cannot be predicted in advance $\cdots$] Development, which is a scheduled activity with a well-defined outcome in a specified time frame, aimed at the marketplace.” (p. 11)
@@
The indirect relationship between theory and practice in the engineering approach to market design is largely shaped by the demands of commissioned projects. Economic designers in such settings lack the flexibility to choose their problems or timelines; they must deliver practical solutions under strict deadlines. As Narayanamurti and Odumosu (2016) note, the pressures of completing a “scheduled” activity often leave little room for advancing theoretical boundaries. This dynamic, however, is not unique to engineering approach to market design—it reflects a broader paradigm articulated by Bush (1945), which envisions a linear, unidirectional relationship between scientific discovery and technological innovation. This model has profoundly influenced U.S. science and technology policy since World War II.
In contrast to the pioneers of market design who were engaged in commissioned projects, I—as an aspiring economic designer—had two significant flexibilities unavailable to them in the late 1990s when I ventured into the world of policy design. I could actively seek out real-world applications with a stronger role for novel theory, and I had ample time to pursue these projects as an “unscheduled quest for new knowledge and the creation of new inventions.”
Freed from these constraints, minimalist market design emerged as an exception to the broader trend shaped by Bush’s linear model. It embraces a holistic approach that integrates research and practice, aligning instead with alternative innovation frameworks advocated by Stokes (1997) and Narayanamurti and Odumosu (2016). To clarify these distinctions, we will first review Bush’s post-war science and technology framework in Section 3.1, then examine Stokes’s Pasteur’s Quadrant model in Section 3.2, and finally discuss the Discovery–Invention Cycles model introduced by Narayanamurti and Odumosu in Section 3.3.
@@ 3.1 Science, the Endless Frontier: Vannevar Bush’s Linear Model of Innovation
Dichotomy Between Basic and Applied Research.
Bush argued that “basic research is performed without thought of practical ends.” He suggested that basic research, aiming to broaden knowledge without concern for practical applications, and applied research, seeking to solve specific practical problems, should remain distinct. According to Bush, focusing on one detracts from the other, as “applied research invariably drives out pure research.”
Linearity of the Four Phases of Innovation.
Bush further asserted that “basic research is the pacemaker of technological progress.” Figure 3 depicts the linear sequence: basic research expands knowledge; applied research creates practical models and innovations; development converts findings into usable outputs; and practical use implements them. Each stage is successive and depends on the previous one.
Figure 3: Bush’s linear innovation model.
@@
The fifty years following Bush’s report witnessed unprecedented research advances, during which the U.S. emerged as a global leader in science. Bush’s linear model of innovation, as well as its influence on shaping institutions like the NSF and other government funding bodies, went largely unchallenged until the mid-1990s. This historical legacy, coupled with the funding structures of U.S. federal agencies like the NSF during that period, contributed to the prominence of Bush’s influential theses in many fields, including economic design.
@@ 3.2 Beyond the Linear Model: Pasteur’s Quadrant
“Science and technology policy would appear to be in a state of crisis. There are many indicators that a crisis does exist in the partnership between the federal government and the American research universities, that the terms of the partnership are increasingly being questioned and re-examined. $\cdots$]
Even if Vannevar Bush is as much a symbol of this period of dramatic progress as its putative architect, it seems fitting that in the 50th anniversary year of the publication of Science: The Endless Frontier, we celebrate the work and the period: reflect on the origins of the Bush paradigm for scientific excellence, take stock of where we currently are in the relationship between science and government, consider whether we are in a period of crisis, and do some serious work on the future shape of the national system of scientific and technological innovation. $\cdots$]
@@
To address the limitations of Bush’s unidimensional approach, Stokes introduces a framework that incorporates a second dimension: the desire to solve practical problems alongside the quest to expand fundamental understanding. As illustrated in Figure 4, his model divides research into three primary quadrants, [^17] There is also a fourth quadrant where research neither extends fundamental understanding nor serves a practical purpose. each named after a key historical figure:
Figure 4: Stokes (1997)’s classification of research programs: the vertical axis measures relevance for advancing knowledge, and the horizontal axis measures relevance for direct use.
@@
Since its introduction in the late 1990s, Stokes’s framework has profoundly influenced innovation policy in the U.S. For example, it inspired the NSF to incorporate a “broader impacts” criterion into its merit review process for grant applications in the early 2010s. [^19] Stokes’s framework has also influenced innovation policy beyond the U.S. For example, since 2011, researchers submitting proposals to the Swiss National Science Foundation have been able to designate their applications as “use-inspired.” In the European Union, Horizon 2020 (2014–2020) and now Horizon Europe (2021–2027) explicitly evaluate projects on both scientific excellence and their “impact pathways.” In the United Kingdom, from 2009 to 2020, applicants to the Economic and Social Research Council and other Research Councils were required to provide a “Pathways to Impact” statement outlining how their work would generate societal benefits. Since 2020, this requirement has been integrated directly into the main research proposal to embed impact planning within the project’s core design. In Canada, beginning in the late 2000s, the Social Sciences and Humanities Research Council, the Natural Sciences and Engineering Research Council, and the Canadian Institutes of Health Research have required applicants to include a “Knowledge Mobilization Plan.” In Australia, the Australian Research Council introduced “Impact and Engagement” criteria in the mid-2010s, which were formalized in the 2018 Engagement and Impact Assessment. In Japan, the Japan Science and Technology Agency has promoted “problem-oriented basic research” as part of the government’s Science and Technology Basic Plans since the early 2000s. Its influence has been so significant that, in 2014, in response to those advocating otherwise, the NSF Assistant Director for Mathematical and Physical Sciences emphasized that the NSF should not entirely abandon support for pure basic research in Bohr’s Quadrant by forcing all basic research into Pasteur’s Quadrant (Crim, 2014):
@@
To better understand how these quadrants manifest in the field of economic design, let us explore specific examples that illustrate research practices within each quadrant. We will focus on seminal works that have not only advanced the field of economic design but have also contributed to their authors receiving the Nobel Memorial Prize in Economic Sciences, thereby underscoring their significance. We begin with Bohr’s Quadrant.